A road bike frame is advertised to weigh no more than 900g. The actual frame weight distribution is normal with standard deviation of 20g. What should the mean weight be so that the advertisement claim is true for at least 99% of all frames?
Answer #1

Solution:-

Given that,

mean = = 900

standard deviation = = 20

Using standard normal table,

P(Z > z) = 99%

= 1 - P(Z < z) = 0.99

= P(Z < z) = 1 - 0.99

= P(Z < z ) = 0.01

= P(Z < -2.33) = 0.01

z =-2.33

Using z-score formula,

x = z * +

x = -2.33* 20+900

x = 853.4